OpenStudy (anonymous):

Calculate the volume of a sample of N2 gas which weighs 0.255 grams if the pressure of the sample is 1.55 atm and the temperature is 37 degrees C.

OpenStudy (anonymous):

I got 0.149 L ...can someone confirm?

OpenStudy (anonymous):

first you need to explicitely write the hipotesis of perfect gas, then you can apply the equation pV=nRT p=1.55atm V=? n=m/M m=0.255*10^(-3)kg and M=2*(14.0067 ± 0.0001) u R=8.3144621(75 J/(mol*K) T=27+273,15 ~= 300K V=0,01465486 m³ = 14,65L

OpenStudy (aaronq):

you used the wrong value for R (needs to be in terms of atm) and you used kg instead of grams when you converted to the mass to moles (molar mass is in grams/mol) \(\sf(1.55~atm)V=\dfrac{0.255~g}{28~g/mol}*(0.08206~\dfrac{L*atm}{mol*K})(300.15~K)\) \(\sf V=\dfrac{\dfrac{0.255~g}{28~g/mol}*(0.08206~\dfrac{L*atm}{mol*K})(300.15~K)}{(1.55~atm)}\) \(\sf V=\dfrac{0.255~g*(0.08206~L*atm)(300.15~K)}{(28~g/mol)(1.55~atm)*{mol*K}}=0.145~\dfrac{L*g*atm*mol*K}{g*atm*mol*K}=0.145~L\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

sorry..